OpenStudy (anonymous):

trigonometry 3) Simplify. a)sec x/tan x b)(sec x – 1)(sec x + 1)

OpenStudy (watchmath):

a) Write \(\sec x\) and \(\tan x\) in terms of \(\sin x\) and \(\cos x\) b) multiply the two and look at the idenity \(\sin^2 x+\cos^2 x=1\)

OpenStudy (michele_laino):

\[\sec x=\frac{ 1 }{ \cos x }\]

OpenStudy (michele_laino):

so: \[\frac{ \sec x }{ \tan x }=\frac{ 1/\cos x }{ \sin x/\cos x }=...\]

OpenStudy (michele_laino):

hint: first fraction times the inverse of the second fraction

OpenStudy (michele_laino):

\[\frac{ 1 }{ \cos x }\frac{ \cos x }{ \sin x }=....\]

OpenStudy (xapproachesinfinity):

the second one is fairly easy just use (a-b)(a+b)=a^2-b^2

OpenStudy (xapproachesinfinity):

then use an identity another form of pythagorean identity

OpenStudy (anonymous):

i still don't get it